Suppose marginal benefit is given by P-9 Q, marginal private cost is given by P-2Q, and marginal external cost 1S 2 What is the socially optimal price? 5. What is the deadweight loss? 6. How much would a corrective (Pigouvian) tax need to be to move the market equilibrium to the socially optimal equilibrium?
Answer #1

olution(s):

1) At equilibrium:

Marginal benefit = Marginal private cost

9 - Q = 2Q

9 = 3Q

Q = 3 units

2. Market price, P = 9 - 3 = $ 6

3. Social optimal:

Marginal benefit = Marginal social cost i.e. MPC + MEC

9 - Q = 2Q + Q

9 = 4Q

Q = 2.25 units

4. Social optimal price = 9 - 2.25 = $ 6.75